1- You deposit $300 each month into an account earning 6% interest compounded monthly.

a) How much will you have in the account in 15 years?

b) How much total money will you put into the account?

2- Suppose you want to have $800,000 for retirement in 30 years. Your account earns 10% interest.

a) How much would you need to deposit in the account each month?

b) How much interest will you earn?


c) How much total interest will you earn?

3- You want to be able to withdraw $50,000 each year for 25 years. Your account earns 5% interest.

a) How much do you need in your account at the beginning?

b) How much total money will you pull out of the account?

c) How much of that money is interest?
Answer #1

Answer:

Question 1)

Deposit each month A=$300

interest rate =6%

Monthly Interest rate r=6%/12=0.5%

a) Amount in account after 15 years P=A*((1+r)^(15*12)-1)/r

P=300*((1+0.5%)^180-1)/0.5%

P=$87245.61

b) Total amount deposited =A*15*12=300*180=$54000